Water is flowing in a fire hose with a speed of 1.0 m/s with a height of 3.0 m and a pressure of 200,000 Pa. You lift the nozzle of the hose to a height of 1.5 m above the ground. At what speed does the water now come out of the hose?

1
Expert's answer
2021-12-09T15:48:09-0500
"200000+0.5(1000)(1)^2+(1000)(9.8)(3)\\=101325+0.5(1000)v^2+(1000)(9.8)(1.5)\\v=15\frac{m}{s}"